Based on past experience, it is assumed that the number of flaws per foot in rolls of grade 2 paper is an average of 1 flaw per 5 feet of paper (0.2 flaw per foot). What is the probability that in a 14 foot roll there will be at least 2 flaws?
Answer #1

Here, λ = 0.2 * 14 = 2.8 and x = 1
As per Poisson's distribution formula P(X = x) = λ^x * e^(-λ)/x!

We need to calculate P(X > 1) = 1 - P(X <= 1).
P(X > 1) = 1 - (2.8^0 * e^-2.8/0!) + (2.8^1 * e^-2.8/1!)
P(X > 1) = 1 - (0.0608 + 0.1703)
P(X > 1) = 1 - 0.2311 = 0.7689